UT 20130912 Technical Observing

Observer: DThompson, JHill, DMiller (Tucson)
Telescope Operator: Sallanson (LBTO)
Software Support: DCox (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: KNewton (LBTO)
Instrument Support: JPower (LBTO)

Summary

Dave T did the first half with LBCs. Data collection for checkout was completed, but there were significant problems with rotator trajectories and RPC errors. Tests were made using LBC internal rotator trajectories.

John H and Doug M did the second half with AGW2. The DX M2 shell has some wrinkles that are visible on the wfs, but imaging performance is identical in 0.5 arcsec seeing. Having proper air pressure to AGW2 fixed our elongated pupil wobble issue, although we still have gotten DX M3 aligned as well as SX M3. We tried to collect field aberration data during the last hour, but conditions (both seeing and clouds) were deteriorating.

Seeing was excellent (0.5 - 0.8 arcsec) for most of the night, although we were interrupted by patches of cloud.

It appears that we need one more night (or better 2 half nights) to finish the AGW2 alignment.

Details

LBCs

Delay in opening because of cloud(s)

OneBias_Bino looks OK, some pickup noise on blue

Open

02:15 Opening

02:30 ACT0302 to do initial rangebal and dofpia

lbcrangebal B=23159 R=23147
   CE  -37.608     1.731   -35.876
   CA -109.506     7.128  -102.378
On LBC-Blue/LBC-Red the star is 15.0/4.1 arcsec away from the rotator center.Based on these, the following mount and optics offsets were calculated to co-point while optimizing the range available to M1: 
MOUNT: delta_IE = 1.73 arcsec, delta_CA = 7.13 arcsec
M1 SX: dX = 0.35 mm, dY = 0.07 mm, dRX = 1.44 arcsec, dRY = -7.56 arcsec 
M1 DX: dX = -0.31 mm, dY = 0.11 mm, dRX = 2.36 arcsec, dRY = 6.68 arcsec 

dofpia, /x2, /first
--------------------------------------------------------------------------
          Computed Aberrations (nm)              |   (arcsec)   | channel 
 Defoc   AstX   AstY   ComaX  ComaY   Sph        |  Est.   Est. |         
 (Z4)    (Z5)   (Z6)   (Z7)   (Z8)   (Z11)  (Z22)| seeing  WIQ  |         
--------------------------------------------------------------------------
  -149     38    -98  -1119   -261     38      0   0.91   0.53     red
  6755      0      0      0      0      0      0   2.01   2.66    blue
   731    207   -180    581     34   -235    -36   0.92   0.43     red
   -87   -288   -249   -159   -181      0      0   1.71   0.18    blue
  -262    -64    185   -143   -274      6     33   0.82   0.20     red
  1746     22    129   -526     17   -602   -100   1.52   0.83    blue
   657   -161    139    561    710   -230    -29   0.82   0.51     red
  -555     72   -342    303    -19    -96      7   1.59   0.29    blue
 -1171    102    200   -608   -962    120      9   0.71   0.70     red
  -784   -260   -231   -307    -41      0      0   1.92   0.36    blue
   407    -82   -128   -539    154    -61      0   0.76   0.31     red
  1307    140    137    119    206   -278    -34   1.54   0.56    blue
    56    293   -400    100    324   -126    -17   0.84   0.24     red
   642     42   -108    372     31   -241    -12   1.61   0.35    blue
  -345     65    117    -77   -619    -40    -10   0.61   0.32     red
  -612     18    162   -138    -89    -49     24   1.60   0.27    blue
    36   -111    -16    115    659    -72    -33   0.64   0.32     red
   131   -108    -17   -123    131     13      1   1.55   0.13    blue
  -145    155    -70    -25    144    -28    -29   0.63   0.13     red
   242    104     31   -172    -19    -90      8   1.55   0.16    blue
Both converged!

ACT0302 in focus...R 25517 B 25530 seeing <1"!

NGC6779 DOFPIA, /X2
   320   -137    116    830   -417    -59     -4   0.51   0.45     red
   597   -113    -82    529     11     35      5   0.92   0.35    blue
   141    136    -89   -407    312    -41     15   0.73   0.26     red
   124     16     87   -206    -23    -37     -4   1.32   0.14    blue
   712    -13    220    -17     68   -207    -19   0.46   0.33     red
   114   -148   -117    111     28    -21    -21   1.19   0.12    blue
 -1294     45    109    230   -207    212    -13   0.54   0.55     red
  -282     63   -129    -81    -24    128     18   1.53   0.17    blue
   889   -161    184     -4    117   -118     17   0.51   0.38     red
   371    125     48   -187    -85     11     22   1.34   0.19    blue
   389   -113    127   -317     -2    -92     11   0.47   0.24     red
   388   -164    -20    -23      2    -46    -27   1.28   0.18    blue
   865    -13    -91    149    108   -185    -13   0.52   0.38     red
  -466    132    -73   -323    199      9     17   1.30   0.27    blue
  -562     25   -195    -73    300      3     13   0.64   0.28     red
   201   -111    100    135    129     21      1   1.11   0.15    blue

Stopped with "S", red not converged but it looks good.

NGC6779_offset_guide.ob
guiding working, rotator problem apparent.

Switch over to using LBC internal trajectories

03:40 updated redchannel.conf/bluechannel.conf on lbccontrol in /home/lbccontrol/conf:
Channel.GetRotatorTrajectoryFromTelescope = FALSE #was TRUE
then lbckill/lbcstart
original versions saved in red/bluechannel.20130912.conf
JMH (the next day): These changes were in LBC software version lbc1.4, and have not yet been propagated to lbc1.5.
KS (later): The images noted in the InformationTechnology using the internal LBC rotator trajectories DID indeed use the internally generated trajectory. There are multiple messages in the log file that indicate this:
2013/09/12 03:45:48.183645 S R                   read ENABLE parameter Channel.GetRotatorTrajectoryFromTelescope< with value FALSE
2013/09/12 03:48:28.169358 W R                   TCS custom code for rotator trajectory [src/tcs/tcs.c:532]
This was the configuration until the instrument was turned off at 06:34UT. LBC used its own calculation of trajectories based on TCS values for TargetDEC, TargetRA, LST.

NGC6779.ob to redo DOFPIA, /X2
  1551    -29     31    467    349    -90     -1   0.44   0.67     red
  -331    -33    -13    331    -61    -26     -7   0.42   0.22     red
  -319     22     30     29   -310    157     -5   0.39   0.23     red
   530     68    -33   -403    395   -181    -18   0.36   0.36     red
  -134    -33    -10     36     70    -27      3   0.38   0.10     red

   247    218    314    569   -330    168     -7   0.86   0.36    blue
  -231      8    208     37     42     22      5   0.86   0.14    blue
  -838    195     25     53     12    282     20   0.76   0.39    blue
  -319   -224   -113     93    123    169      0   0.93   0.21    blue
   127     33    337     56    -60    -22     -3   1.01   0.14    blue
both converged

Note to Michelle, these globular cluster fields are a bit dense, DOFPIA is having trouble fitting the sky. Move off another ~10 arcmin?

03:59 NGC6779_offset_guide.ob
lbcb.20130912.035927, lbcr.20130912.035923
blue guidestar on postage stamp on GUI is elongated, red looks OK
red looks good, ~0.7", blue is elongated 2:1
no apparent rotator problems
OB stopped after second exposures read out.

04:14 NGC6779.ob for DOFPIA
  -662     30     40    281    -98      1      9   0.46   0.30     red
   226    -23    -39   -104    125    -72      0   0.46   0.15     red

  -275   -125   -159     43   -293   -275    -16   0.86   0.26    blue
   111    121    -50     22      6   -102    -11   0.75   0.12    blue

Focal Plane Tilt
04:19 RB_rV_superfoc7 (PA=0)
  • B=042004 R=041958
  • 042054 042037
  • 042143 042117
  • 042232 042158
  • 042321 042238
  • 042410 042318
  • 042500 042359

04:26 RB_rV_superfoc7_pa180
  • B=042750 R=042745
  • 042839 042825
  • 042929 042905
  • 043018 (no more reds)
  • 043107
  • 043155
  • 043245

Red error:
112   2013/09/12   04:30:39.081750         OB execution error: one or more systems failed [src/lbc/actions.c:1159]
111   2013/09/12   04:30:38.941805         >>> preset/expose/save interrupted at expose time 
RPC error, restart OB gave another RPC error.  

Turn off, turn on

93   2013/09/12   04:34:34.509004         camera system encountered a problem 6 [src/channel/channel.c:680]
92   2013/09/12   04:34:24.040401   CAMERA       RPC Client/Host communication failed retval:3 [src/camera/camera.c:643]
91   2013/09/12   04:33:30.676917         OB execution error: one or more systems failed [src/lbc/actions.c:1159]
90   2013/09/12   04:33:30.546853         >>> preset/expose/save interrupted at preset time
89   2013/09/12   04:33:28.726773         camera system encountered a problem 6 [src/channel/channel.c:1104]
88   2013/09/12   04:33:28.726659   CAMERA       RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101] 

04:40 restarted RB_rV_superfoc7_pa180
  • B= R=044037 (blue never started)
  • abort and restart

04:43 restarted RB_rV_superfoc7_pa180
  • B=044320 R=044306
  • 044410 044347
  • 044500 044426
  • 044549 044506
  • 044638 044546
  • 044727 044626
  • 044819 044706

04:47 DOFIPA, /X2
  -794    -15      9    497     45    113     19   0.40   0.40     red
   -52     10     26   -138     82      1      1   0.44   0.11     red

  -599    158     54     19     18    -83      7   0.72   0.26    blue
   146   -132    -72     57     25    -53     -1   0.83   0.12    blue

RB_rV_superfoc7_pa180
  • B=045526 R=045520
  • 045616 045559
  • 045705 045639
  • 045754 045719
  • 045843 045759
  • 045933 045839
  • 050022 045919

RB_rV_superfoc7 (PA=0)
  • B=050329 R=050322
  • 050419 050402
  • 050507 050442
  • 050556 050522
  • 050645 050602
  • 050735 050641
  • 050824 050721

DOFIPA, /X2 (EL=65)
   186     46     30    -41   -107    -28      6   0.43   0.12     red
  -145     -9     37   -183    109     15      9   0.46   0.14     red
   130     53     51    278   -142    -49      1   0.42   0.17     red
   345    -13     27   -333     58   -145      2   0.40   0.24     red

  -371    105     19      9     50   -193    -12   0.80   0.21    blue
   271    182    -93     12    -92   -103     -6   0.74   0.17    blue
  -313   -132    -23    -93      0     -3      0   0.69   0.16    blue
    -3   -118     72   -108     -7    -67     -4   0.69   0.11    blue

05:15 Filter Focus Offsets: g-SLOAN
  • RB_rV_superfoc7.ob B=51701 R=51655 @ start
  • RB_ig_superfoc7.ob B=52335 R=52314

Another red RPC error
   2013/09/12   05:25:24.451010         OB execution error: one or more systems failed [src/lbc/actions.c:1159]
2   2013/09/12   05:25:24.371026         >>> preset/expose/save interrupted at preset time
1   2013/09/12   05:25:24.370921         camera system encountered a problem 6 [src/channel/channel.c:1104]
0   2013/09/12   05:25:24.370801   CAMERA       RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101] 

telescope@lbccontrol appears not to have permission to kill/start LBC:
lbto@obs2:22 % ssh telescope@lbccontrol
telescope@lbccontrol's password: 
[telescope@CMU telescope]$ cd /home/lbccontrol
[telescope@CMU lbccontrol]$ ./lbckill.sh 
lbc(30339): Operation not permitted
lbc(30344): Operation not permitted
lbc: no process killed
[telescope@CMU lbccontrol]$ exit
Done as root which worked

05:30 DOFIPA, /X2 (EL=60)
  -448    -10     17    484    -12      0      2   0.45   0.29     red
    40     52     -9   -315    155    -84     -5   0.44   0.19     red

  -509   -113    -93    201    -42    -97      1   0.79   0.25    blue
   181    111    -24    -81     88     -5     -3   0.74   0.12    blue

05:36 Filter Focus Offsets: g-SLOAN (again)
  • RB_rV_superfoc7.ob
  • RB_ig_superfoc7.ob
  • RB_rV_superfoc7.ob red warning again

9   2013/09/12   05:51:29.808118         OB execution error: one or more systems failed [src/lbc/actions.c:1159]
8   2013/09/12   05:51:29.728077         >>> preset/expose/save interrupted at preset time
7   2013/09/12   05:51:29.727974         camera system encountered a problem 6 [src/channel/channel.c:1104]
6   2013/09/12   05:51:29.727866   CAMERA       RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101] 

blue completed so there is a full set of superfoc for the g filter.

11:02 WT10_306
DOFPIA, /X2

  3209    -28   -108   -526   -402   -152     -6   0.62   1.31     red
   237    -11    -73    290   -147   -156     -1   0.53   0.22     red
  -198    -32    -28     -5    -17    -80     18   0.54   0.12     red
   121     22    -19     52      6    -88     11   0.57   0.11     red

  2094   -268   -103   -389   -199    -95     10   0.89   0.86    blue
   682   -176     30    -36    151   -218     -6   0.86   0.32    blue
   379    -73    -48    152    124   -218     -7   0.81   0.24    blue
   163     76    -63   -174   -101    -65     15   0.89   0.15    blue

Run guide test again, 360x on WT10_306
Rotator problem persists! EL>85

06:25 Turning telescope over to Doug Miller for AGw#2 tests. Reconfiguring to FrontBent on both sides

IRTC@RFBG with AGW2

07:00 Preset fails when image doesn't arrive from AzCam. Restarting GCSR didn't help. Steve is rebooting various AzCam computers.

When we were last here on the morning of 20130903 UT, the RFBG pupil wobble had converged to a line rather than a point. This has been fixed by putting the compressed air supply for AGW2 back to its nominal pressure.

07:15 Steve has cold booted the azcamserver (5.1) machine, but now it doesm't respond to ping. ??

07:20 Call Taras. He says we should be using azcamserver (4.6).

07:26 ACQUIRE preset succeeds, but still no AzCam image. Call Taras again. We switched back to the "new box".

07:36 Authorize LUCI1 and IRTC

07:40 ACQUIRE preset with BC=0 -- succeeds with a movie.

07:45 ACTIVE preset to BS9101 BC=0 -- Acquisition image looks like the space shuttle. Seeing is 0.55 - 0.7 on the guider at EL=64.

07:55 async ACTIVE preset on left - some binocular tests while Doug is tuning his script

07:57 async ACQUIRE - Steve picks the correct focus direction. Needed 0.4 mm.

07:59 ACTIVE preset fails on left when acquisition picks the edge of the bottom black stripe (lower right corner).

08:06 Clouds

08:13 The good news about these clouds is that the guider is no longer grabbing the hot pixels when the star disappears. The GCS patch 2.12.6 is working correctly, but is numbered on the GUI as 2.12.5.

Pupil wobble with AGw2 @ RFBG

/home/lbto/data/20130912/sh_pupil_wobble/set_?

08:20 Doug is collecting pupil wobble data on the right side. Tip=-200, Tilt=+250
pupil_collect - set_1

08:35 The thicker clouds are gone, and seeing is bouncing around in the 0.4 to 0.6 arcsec range on the guiders.

08:47 Move DX M3 Global Offset Tip from -200 to 0 in steps of 50. Tilt remains at +250

08:50 Move DX M3 Global Offset Tilt from +250 to 0 in steps of 50. Tip remains at 0

08:57 Move DX M3 Global Offset Tilt from 0 to -250 in steps of 50. Tip remains at 0

These steps of 50/5 seem to give 1 arcsec jumps on SX.

09:01 Move DX M3 Global Offset Tilt from -250 to -500 in steps of 50. Tip remains at 0

09:07 Move DX M3 Global Offset Tip from 0 to +200 in steps of 50. Tilt remains at -500

09:10 pupil_collect - set_2

guide.telescope.tmp is colliding between sx and dx for two instances of LBTplot.

09:25:27 stop reporting centroids on GCSL

09:30:57 resume reporting centroids on GCSL

09:35 Move DX M3 Global Offset Tip from +200 to 0 in steps of 50. Tilt remains at -500

09:39 Move DX M3 Global Offset Tilt from -500 to -250 in steps of 50. Tip remains at 0

09:43 Move DX M3 Global Offset Tilt from -250 to -150 in steps of 50. Tip remains at 0

09:46 pupil_collect - set_3

10:12 Move DX M3 Global Offset Tilt from -150 to 0 in steps of 50. Tip remains at 0

10:xx pupil_collect - set_4 --> radius is about 4 pixels.

10:15 Seeing "explodes" from 0.5 to 1.2 arcsec as some more clouds arrive.

--> radius is about 4 pixels. Why can't we make it smaller? Is something else out of alignment?

Field data for AGw2 at BS9107

/home/lbto/data/20130912/field_agw2/

10:42 Authorize IRTC @ RFBG only

10:43 ACQUIRE preset to recover pointing since we moved M3 IE=-2 CA=-133

10:50 ACTIVE preset to BS9107 at EL=81

10:58 field_collect with BS9107_new_field.list

But the on-axis pupils are not well centered.

11:10 Doug measured a new on-axis pupil center at 371.5, 253.8 ---> John applied this to WFSCam_hotspot_x in LUCI_IRTC_R.cfg (takes effect without restarting GCSR)

11:15 field_collect with BS9107_new_field.list - seeing on the guider is running 0.7 to 0.85 arcsec FWHM at EL=85.

11:27 Having trouble collimating below 250 nm on-axis. Seeing doesn't seem that bad, so I'm suspecting we got some stray Z9 while the pupil position was off.

11:30 clear active optics on DX M1

11:32 field_collect, start=2 with BS9107_new_field.list - seeing on the guider is running 0.7 to 0.85 arcsec FWHM at EL=85.

starting with image 85 convergence is much better - don't use the data above.........

GStar wfsc #
0 85-86
28 87-89
0 90-91
36 92-94
0 95-97
33 98-100
0 104-103
18 106-108
GStar wfsc #
0 109
19 110-112
0 116-118
24 119-121
0 122-124
42 129-131
0 132-134
20 135-137
0 138-140
50 141-143

11:44 clouds are covering a larger fraction of the sky - including the part we are looking through.

There is alot of vignetting in wfsc 00098 on gs=33

This data isn't very good, but it is clear that we've got some binodal on DX M2 that needs adjusted out.

11:55 missed the star in acq

11:56 ACQUIRE preset IE=+11 CA=-128

11:57 field_collect, start=6 with BS9107_new_field.list

12:25 tadpole images show that hotspot position is OK, but higher order aberrations make it difficult to make a precise determination. Maybe the star is a bit to the left in the field stop.

12:29 IDL printed aberration values from the previous image set. ??

12:30 on-axis preset lost in clouds

Results from analysis of field data 20130912 DMiller

 => Field Data Results
             Measured bi-nodal astigmatism
  
                  rfbg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  37.7", Delta IE = -37.7"
         Tilt = 127.4", Delta CA = -127.4"
  
  
                  rfbg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  30.6", Delta IE = -30.6"
         Tilt = 122.8", Delta CA = -122.8"

* 45-Astigmatism plane:
45_astig.jpg

  • Focus plane:
    focus.jpg

Close

12:33 close for impending dawn

-- DavidThompson - 12 Sep 2013
Topic attachments
I Attachment Action Size Date Who Comment
45_astig.jpgjpg 45_astig.jpg manage 66 K 13 Sep 2013 - 03:44 DougMiller 45-Astigmatism plane
focus.jpgjpg focus.jpg manage 69 K 13 Sep 2013 - 03:45 DougMiller Focus plane
Topic revision: r9 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback